OpenStudy (anonymous):

What are the first five terms of a1=7 d=-5

OpenStudy (cwrw238):

this is an arithmetic series a1 = 7 the other terms will be 7 + (n-1)d so second term is 7 +(2-1)* -5 = 7 -5 = 2 can you continue ?

OpenStudy (cwrw238):

n = 2nd 3rd 4th term etc

OpenStudy (anonymous):

my options are -5,2,9,16,23 7,2,-3,-8,-13 7,12,17,22,27 2,9,16,23,30

OpenStudy (cwrw238):

well as you can see i have done first 2 they are 7,2

OpenStudy (cwrw238):

these others are all going down by 5

OpenStudy (anonymous):

gotcha! that was pretty dumb i totally missed that haha sorry! Thanks!

OpenStudy (cwrw238):

yw

Parth (parthkohli):

There's an easier approach to this tho

OpenStudy (anonymous):

oh?

Parth (parthkohli):

The common difference is -5 so you are again and again adding -5. Just keep subtracting 5 again and again to get the next terms :)

Parth (parthkohli):

7 - given 2nd term => 7 - 5 = 2 3rd => 2 - 5 = -3 etc.

OpenStudy (anonymous):

yeah i must have punched that in my calculator wrong cause i wasn't getting that before haha i am getting it now